OpenStudy (anonymous):

CAN ANYONE SOLVE THIS?? 2^x = 10

OpenStudy (anonymous):

x=3.33

OpenStudy (anonymous):

r ealyy

OpenStudy (anonymous):

2^3.33=10.0

OpenStudy (anonymous):

can you also solve this one ?? 3^(1 – 2x) = 4

OpenStudy (anonymous):

u still there @Gikundab ??

OpenStudy (anonymous):

x=1.27

OpenStudy (anonymous):

do u mind solving a few more for me??

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2^(x + 1) = 5^(1 – 2x)

OpenStudy (anonymous):

actually x is 2.27 because 3^(1-2.27)=4 3^1.27=4.0

OpenStudy (anonymous):

can you solve this on e 2^(x + 1) = 5^(1 – 2x)

OpenStudy (anonymous):

u there??

OpenStudy (anonymous):

(x+1).log2 = (1-2x).log5 0.30103x + 0.30103 = 0.699 - 1.398x (0.30103+1.398)x = 0.699 - 0.30103 1.699x = 0.398 x = 0.398/1.699 = 0.2343 >================< ANSWER

OpenStudy (anonymous):

what about this one ?? 5(2x) = 36

OpenStudy (anonymous):

oh hold on i did the wrong one its x=3.6

OpenStudy (anonymous):

need and explanation?

OpenStudy (anonymous):

i got x=2.84

OpenStudy (anonymous):

2^2.84=7.2

OpenStudy (anonymous):

5(2x)=36 10x=36 10x/10=36/10 x=3.6